OpenStudy (anonymous):

when you do fractions does the biggest number go on top if you Simplify like say 5/12 divided 11/12 11/5 is the answer

OpenStudy (anonymous):

(5/12)/(11/12)=(5*12)/(12*11)=5/11

OpenStudy (anonymous):

So really the other way around

OpenStudy (anonymous):

man i did have it right and i flipped it thank you

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

the biggest number doesn't necessarily go on top (5/12)/(10/3) = (5/12) * (3/10) you take the reciprocal of the denominator and multiply